Following table shows demand of product at different prices.
Price (Rs.) Demand (Units)
1000 5000
900 6000
825 6800
700 7500
550 8500
a. Draw the demand curve for the above table.
b. If the store decided to increase the price of its desktop computers from
$825 to $900, the stores sales will fall by how many computers?
c. Find out the elasticity at each demand point.
d. Is the price elasticity of demand elastic, unit elastic, or inelastic?
e. Will the store’s total revenue increase, decrease, or remain unchanged?
1
Expert's answer
2021-10-04T10:18:22-0400

a.




b

"6800-6000=800"


c.

"\frac{1000}{-100}\times\frac{1000}{5000}=-2"


"\frac{800}{-75}\times \frac{900}{6000}=-1.6"


"\frac{700}{-125}\times \frac{825}{6800}-0.68"


"\frac{1000}{-250}\times\frac{700}{7500}=-0.37"


d.

price elasticity is elastic


e

The total revenue increases